User Tools

Site Tools


handover:aua012

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Last revision Both sides next revision
handover:aua012 [2016/08/09 18:10]
Lucas Hyland
handover:aua012 [2016/08/10 00:01]
Jamie McCallum
Line 7: Line 7:
  
    * 17:30UT: Experiment Started OK (Tiege)    * 17:30UT: Experiment Started OK (Tiege)
 +   * approx 2300 - Power outage at Katherine. Antenna drives off, rfpic server needed restarting. GPIB converter powered off and unable to be power cycled. Some observations made in this period, but without clkoff measurements. (JMc) 
 +   * 2337 -  DBBC rebooted itself - power glitch?. Martin arrived at station, fixed GPIB error. Schedule resumed with first good data at 2356. (JMc) 
 +   * All recorded data between 2300 and 2355 is suspect. (JMc)
  
 **Yarragadee** **Yarragadee**
Line 21: Line 23:
    * 17:30UT: Experiment did not start, major problems with the set up. setupsx was getting the critical error (sample rate error as stated in the setup guide), editing the proc file and making it similar too aua011 got rid of this error but introduced others. On-call person was called (Warren) we spent half an hour mucking around with the proc file, Warren eventually hacked together a proc file that seems to work without spitting out errors, however, neither of us are sure it is working correctly. Will wait for Jamie to evaluate at shift change over. Warren edited something to do with the tracks in the proc file. (Tiege)    * 17:30UT: Experiment did not start, major problems with the set up. setupsx was getting the critical error (sample rate error as stated in the setup guide), editing the proc file and making it similar too aua011 got rid of this error but introduced others. On-call person was called (Warren) we spent half an hour mucking around with the proc file, Warren eventually hacked together a proc file that seems to work without spitting out errors, however, neither of us are sure it is working correctly. Will wait for Jamie to evaluate at shift change over. Warren edited something to do with the tracks in the proc file. (Tiege)
    * 17:53UT: First good scan 222-1753 (Tiege)    * 17:53UT: First good scan 222-1753 (Tiege)
 +   * 2300: Configuration confirmed as valid as far as frequencies and fanout. Will need to send a revu=ised VEX file to SHAO for the correlation. (JMc)
 +   * 0000: Track assignment may be problematic.
  
  
/home/www/auscope/opswiki/data/pages/handover/aua012.txt · Last modified: 2016/08/10 00:16 by Jamie McCallum